Question 1

An airline company wishes to know the proportion of business class travelers flying the Kuala Lumpur-to-Hong Kong route. In a random sample of 350 passengers, 190 are business class passengers.

 

a. Determine the point estimate of the true proportion of business class passengers.

b. Construct a 95% confidence interval estimate of the average waiting time for all customers.

c. Construct a 99% confidence interval estimate of the average waiting time for all customers.

d. Referring to b) and c), what will happen to the width of confidence interval?



1
Expert's answer
2021-04-15T06:30:38-0400

a. The point estimate of the true proportion of business class passengers

"\hat{p}=\dfrac{X}{N}=\dfrac{190}{350}=\dfrac{19}{35}\approx0.542857"



b. The critical value for "\alpha=0.05" is "z_c=z_{1-\alpha/2}=1.96."

The corresponding confidence interval is computed as shown below:


"CI(proportion)"

"=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{19}{35}-1.96\sqrt{\dfrac{\dfrac{19}{35}(1-\dfrac{19}{35})}{350}},"

"\dfrac{19}{35}+1.96\sqrt{\dfrac{\dfrac{19}{35}(1+\dfrac{19}{35})}{350}})"


"=(0.491,0.595)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.491<x<0.595," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.491,0.595)."


c. The critical value for "\alpha=0.01" is "z_c=z_{1-\alpha/2}=2.576."

The corresponding confidence interval is computed as shown below:


"CI(proportion)"

"=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{19}{35}-2.576\sqrt{\dfrac{\dfrac{19}{35}(1-\dfrac{19}{35})}{350}},"




"\dfrac{19}{35}+2.576\sqrt{\dfrac{\dfrac{19}{35}(1+\dfrac{19}{35})}{350}})"


"=(0.474,0.611)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.474<x<0.611," which indicates that we are 99% confident that the true population proportion "p" is contained by the interval "(0.474,0.611)."


d. If you want a higher level of confidence, that confidence interval will be wider.

Increasing the confidence will increase the margin of error resulting in a wider interval.